OpenStudy (pythagoras123):

\log _7\left(3\right)=0.323

OpenStudy (unklerhaukus):

`\[\log _7\left(3\right)=0.323\]` \[\log _7\left(3\right)=0.323\]

OpenStudy (e.mccormick):

You do not really need the left and right tags for a number. (3) works fine alone: `\(\left(3\right)\)` = \(\left(3\right)\) `\((3)\)` = \((3)\) Where those are useful is for fractions and other larger items: `\(\left(\dfrac{3}{x}\right)\)` = \(\left(\dfrac{3}{x} \right)\) `\((\dfrac{3}{x})\)` = \((\dfrac{3}{x})\)